# Can LLMs Effectively Leverage Graph Structural Information through Prompts, and Why?

### v3

@@ Abstract
Large language models (LLMs) are gaining increasing attention for their capability to process graphs with rich text attributes, especially in a zero-shot fashion. Recent studies demonstrate that LLMs obtain decent text classification performance on common text-rich graph benchmarks, and the performance can be improved by appending encoded structural information as natural languages into prompts. We aim to understand why the incorporation of structural information inherent in graph data can improve the prediction performance of LLMs. First, we rule out the concern of data leakage by curating a novel leakage-free dataset and conducting a comparative analysis alongside a previously widely-used dataset. Second, as past work usually encodes the ego-graph by describing the graph structure in natural language, we ask the question: do LLMs understand the graph structure in accordance with the intent of the prompt designers? Third, we investigate why LLMs can improve their performance after incorporating structural information. Our exploration of these questions reveals that (i) there is no substantial evidence that the performance of LLMs is significantly attributed to data leakage; (ii) instead of understanding prompts as graph structures as intended by the prompt designers, LLMs tend to process prompts more as contextual paragraphs and (iii) the most efficient elements of the local neighborhood included in the prompt are phrases that are pertinent to the node label, rather than the graph structure  [^1] Codes and datasets are at: https://github.com/TRAIS-Lab/LLM-Structured-Data.
@@ 1 Introduction
To this end, this paper focuses on three concrete questions relevant to the incorporation of structural information into LLMs. First, we seek to understand whether the performance gain of LLMs on common node classification datasets comes from data leakage. Second, we investigate whether LLMs understand the graph structure as intended by the prompt designer. Third, we investigate potential reasons why LLMs benefit from structural information.
@@
Our key findings are summarized as follows. (i) There is no strong evidence that data leakage is a major factor contributing to the performance of LLMs on node classification benchmark datasets. (ii) LLMs understand the prompts as linearized paragraphs, despite the prompt designers’ intent to describe the graph structure in natural language. (iii) The most efficient elements of the local neighborhood included in the prompt are phrases that are pertinent to the node label, rather than the graph structure.
@@ 2 Related Literature
Data leakage in LLMs has become a focal point of discussion due to the models’ intrinsic ability to memorize training data. As demonstrated by Carlini et al. (2022), LLMs can emit memorized portions of their training data when appropriately prompted, a phenomenon that intensifies with increased model capacity and training data duplication. While memorization is inherent to their function, it raises serious security and privacy concerns. A study by Carlini et al. (2021) shows that extraction attacks can recover sensitive information such as personally identifiable information (PII) from GPT-2 (Radford et al., 2019). This capability to store and potentially leak personal data is further explored by Huang et al. (2022), confirming that although the risk is relatively low, there is a tangible potential for information leakage. Specifically,  Carlini et al. (2022) show that the 6 billion parameter GPT-J model (Wang & Komatsuzaki, 2021) memorizes at least 1% of its training dataset. Furthermore, the issue of data leakage complicates the evaluation of these models. As highlighted by Aiyappa et al. (2023), the closed nature and continuous updates of models like ChatGPT make it challenging to prevent data contamination, affecting the reliability of evaluation on LLMs in various applications. In node classification tasks, a concurrent work by Chen et al. (2023) observe that a specific prompt alteration significantly improved performance on ogbn-arxiv, raising concerns about potential test data leakage. In this work, we take a rigorours approach by curating a new dataset for node classification tasks, which is explicitly designed to address the data leakage issues in existing benchmarks.
@@
Recently there has been a series of research on LLMs and text-rich graphs. He et al. (2023) propose a method where LLMs perform zero-shot predictions along with generating explanations for their decisions, which are then used to enhance node features for training Message Passing Neural Networks (MPNNs) (Gilmer et al., 2017) to predict node categories. Chen et al. (2023) extend the work of  He et al. (2023) by using LLMs both as feature enhancers and as predictors for node classification. They offer several observations such as Chain-of-thoughts is not contributing to performance gains. Guo et al. (2023) perform an empirical study on using LLMs to solve structure and semantic understanding tasks. More recently, Ye et al. (2023) propose InstructGLM for the instruction tuning of LLMs, like LLaMA (Touvron et al., 2023), for node classification tasks.  Zhao et al. (2023) introduces= GraphText to encode graph into natural languages using a graph-syntax tree. Our study differs with this line of research in terms of the motivation: while we are using text-rich graph datasets as a testbed, our primary goal is to gain deeper understanding of LLMs’ capability of processing the graph modality, instead of leveraging LLMs to better solve node classification tasks.
@@ 3.1 Research Questions
In this section, we aim to gain a deeper understanding of three central questions. First, does the classification performance of LLMs on common node classification datasets come from data leakage? Second, do LLMs understand the graph structure as intended by the prompt designers? Third, what factors contribute to LLMs benefiting from structural information? To ground our study, we experiment with the ChatGPT API and LLaMA-2-7B model on node classification datasets that have textual node features.
@@ 3.2 Data Leakage as a Potential Contributor of Performance
While LLMs have achieved decent classification performance on common node classification datasets, there is a risk that the performance of LLMs is artificially inflated by data leakage (He et al., 2023; Guo et al., 2023; Hu et al., 2023). Note that most node classification benchmark datasets have a data cut-off at 2019 (see Table 9 in Appendix B.1), but ChatGPT was trained on data up to September 2021 (OpenAI, 2023) and LLaMA-2 was trained on data up to September 2022  [^2] https://github.com/facebookresearch/llama/blob/main/MODEL_CARD.md. While the training datasets of ChatGPT and LLaMA-2 are not publicly available, given the widespread of these datasets on the internet and the enormous training corpus of ChatGPT and LLaMA-2, it is reasonable to worry about the data leakage issue on these datasets.
@@
3.3 Are LLMs Treating the Prompts as Graphs as Intended by Prompt Designers?
Previous work designs templates to encode the ego-graph around the target node into natural languages (Chen et al., 2023; Guo et al., 2023; Hu et al., 2023). One major similarity of these templates is sequentially describing the graph structure in natural language, starting with the immediate (1-hop) neighbors of the target node and then progressing to the subsequent levels (2-hop, 3-hop, etc.). However, it is unclear whether LLMs actually understand the prompt as intended. Thus, we ask the question: do LLMs treat the input prompt as a graph, as expected by the prompt designers, or merely as a paragraph of text with augmented keywords? This question is fundamental in understanding the performance gain after incorporating structural information. If LLMs just process the prompt as a paragraph consisting of neighboring papers’ titles, then the performance gain can only come from the extra phrases included in neighboring nodes, instead of the topological information. We investigate this question by two adversarial experiments:
@@ 3.3 Are LLMs Treating the Prompts as Graphs as Intended by Prompt Designers?
To verify whether prompts are processed as linearized paragraphs, we create a linearized version of $k$-hop title, $k$-hop title+label and $k$-hop attention by stripping away descriptive text about the ego-graph structure from the original prompt. For instance, in the original $k$-hop title, a typical introduction like “It has the following neighbor papers at hop $k$:” is removed. In the linearized version, we condense this to a single paragraph comprising all neighbors’ titles, as shown in the last row of Table 1. We then compare the performance of the graph-structured and linearized prompts on ChatGPT. As indicated in Table 4, the results reveal minimal performance disparity across all five prompt styles after linearization  [^6] The model used here is ChatGPT in September, 2023. A notable finding is that in the $1$-hop attention category, even the elimination of text specifying neighbors’ importance ranking (“It has the following important neighbors, from most related to least related”) has little impact. These findings imply that LLMs tend to process prompts in a linearized paragraph format.
@@
In summary, our study suggests that LLMs interpret inputs more as contextual paragraphs than as graphs with topological structures, which contradicts the intention of prompt designers. By two adversarial experiments, we show that neither linearizing nor rewiring ego-graph has significant impact on the classification performance of LLMs.
@@ 3.4 Impact of Homophily on LLMs Classification Accuracy
We design a controlled experiment to demonstrate the effect of local homophily ratio on prediction accuracy. We gradually drop neighbors in three different ways: a) drop the neighbors with same label as the target node; b) drop the neighbors with different label as the target node; and c) drop neighbors randomly. We include details about the neighbor dropping strategies in Appendix E. The experimental results are shown in Figure 3, where we observe an evident trend: as we selectively remove neighbors sharing the same labels, there’s a decrease in prediction accuracy. Conversely, discarding neighbors with different labels leads to an increase in accuracy. This selective dropping inherently modifies the local homophily ratio within the prompts. The results show that accuracy of predictions made by LLMs is positively related to local homophily ratio. Combine this observation with the premise that LLMs actually understand the prompts as a paragraph, we propose that the significance of homophily is intrinsic. This is because LLMs benefit from structural information only when the neighborhood is homophilous, meaning the neighbors contain phrases related to the groundtruth label of the target node. Conversely, diminishing the local homophily ratio by excluding such neighbors can adversely affect prediction accuracy.
@@
For the other datasets, a positive correlation is evident in both the zero-shot and 1-hop title+label settings. In Table 6, zero-shot prediction (the one that doesn’t use structural information at all) also showed high correlation with the homophily ratio of the node. This suggests a complicated mechanism for LLMs to perform better on homophilous nodes: those nodes are easier to be classified in the first place; the added structural information has some additional contributions.
@@ 4 Conclusions and Future Work
In conclusion, our study provides key insights into the application of LLMs in processing structured data, particularly in node classification tasks on text-rich graphs. In this study, we have adapted node classification datasets with textual features from graph learning benchmarks to establish a testbed for LLMs augmented with structured data. By curating a new dataset, we show that data leakage is not the major contributor of LLMs on common node classification benchmarks. Our findings also raise critical concerns regarding the actual depth of understanding that LLMs have of structured data. By two adversarial experiments, we show that LLMs actually do not understand the prompts as graphs, which is not intended by prompt designers. As a consequence, LLMs only show improvement from structural information when the neighborhood is homophilous. Moreover, our findings suggest that when the target node itself contains a wealth of relevant phrases, the additional structural information becomes redundant.
@@ Appendix A Details about Prompting Format and Settings
ogbn-arxiv, arxiv-2023
Please predict the most appropriate arXiv Computer Science (CS) sub-category for the paper. The predicted sub-category should be in the format ’cs.XX’.
cora
@@ D.1 Classification accuracy on LLaMA-2-7B-chat
Appendix E Additional Analysis for Data Leakage
@@
System Prompt
Zero-shot 1-hop title+label
@@
Please predict the most appropriate arXiv Computer Science (CS) sub-category for the paper. The predicted sub-category should be in the format ’cs.XX’.
74.0 73.7 74.3 70.4
Please predict the most appropriate arXiv Computer Science (CS) sub-category for the paper. Your answer should be chosen from cs.AI, ..cs.SY. The predicted sub-category should be in the format ’cs.XX’.
66.0 68.1 70.7 67.9
Please predict the most appropriate original arXiv identifier for the paper. The predicted arxiv identifier should be in the format ’arxiv cs.xx’.
71.3 70.8 73.7 67.5
Please predict the most appropriate original arXiv identifier for the paper. Your answer should be chosen from cs.ai,.. cs.sy. The predicted arxiv identifier should be in the format ’arxiv cs.xx’.
58.4 57.2 71.7 64.2
Please predict the most appropriate category for the paper. Your answer should be chosen from "Artificial Intelligence",.. "Systems and Control".
54.6 53.4 74.1 67.8

### v4

# Can LLMs Effectively Leverage Graph Structural Information through Prompts in Text-Attributed Graphs, and Why?

@@ Abstract
Large language models (LLMs) are gaining increasing attention for their capability to process graphs with rich text attributes, especially in a zero-shot fashion. Recent studies demonstrate that LLMs obtain decent text classification performance on common text-rich graph benchmarks, and the performance can be improved by appending encoded structural information as natural languages into prompts. We aim to understand why the incorporation of structural information inherent in graph data can improve the prediction performance of LLMs. First, we rule out the concern of data leakage by curating a novel leakage-free dataset and conducting a comparative analysis alongside a previously widely-used dataset. Second, as past work usually encodes the ego-graph by describing the graph structure in natural language, we ask the question: do LLMs understand the prompts in graph structures? Third, we investigate why LLMs can improve their performance after incorporating structural information. Our exploration of these questions reveals that (i) there is no substantial evidence that the performance of LLMs is significantly attributed to data leakage; (ii) instead of understanding prompts as graph structures, LLMs tend to process prompts more as contextual paragraphs and (iii) the most efficient elements of the local neighborhood included in the prompt are phrases that are pertinent to the node label, rather than the graph structure [^1] Codes and datasets are at: https://github.com/TRAIS-Lab/LLM-Structured-Data.
@@ 1 Introduction
To this end, this paper focuses on three concrete questions relevant to the incorporation of structural information into LLMs. First, we seek to understand whether the performance gain of LLMs on common node classification datasets comes from data leakage. Second, we investigate whether LLMs understand the graph structure as topological graphs or contextual paragraphs. Third, we investigate potential reasons why LLMs benefit from structural information.
@@
Our key findings are summarized as follows. (i) There is no strong evidence that data leakage is a major factor contributing to the performance of LLMs on node classification benchmark datasets. (ii) LLMs understand the prompts as linearized paragraphs,rather than as explicit graph structures. (iii) The most efficient elements of the local neighborhood included in the prompt are phrases that are pertinent to the node label, rather than the graph structure.
@@ 2 Related Literature
Data leakage in LLMs has become a focal point of discussion due to the models’ intrinsic ability to memorize training data. As demonstrated by Carlini et al. (2022), LLMs can emit memorized portions of their training data when appropriately prompted, a phenomenon that intensifies with increased model capacity and training data duplication. While memorization is inherent to their function, it raises serious security and privacy concerns. A study by Carlini et al. (2021) shows that extraction attacks can recover sensitive information such as personally identifiable information (PII) from GPT-2 (Radford et al., 2019). This capability to store and potentially leak personal data is further explored by Huang et al. (2022), confirming that although the risk is relatively low, there is a tangible potential for information leakage. Specifically,  Carlini et al. (2022) show that the 6 billion parameter GPT-J model (Wang & Komatsuzaki, 2021) memorizes at least 1% of its training dataset. Furthermore, the issue of data leakage complicates the evaluation of these models. As highlighted by Aiyappa et al. (2023), the closed nature and continuous updates of models like ChatGPT make it challenging to prevent data contamination, affecting the reliability of evaluation on LLMs in various applications. In node classification tasks, a concurrent work by Chen et al. (2023) observe that a specific prompt alteration significantly improved performance on ogbn-arxiv, raising concerns about potential test data leakage. In this work, we take a rigorous approach by curating a new dataset for node classification tasks, which is explicitly designed to address the data leakage issues in existing benchmarks.
@@
Recently there has been a series of research on LLMs and text-rich graphs. He et al. (2023) propose a method where LLMs perform zero-shot predictions along with generating explanations for their decisions, which are then used to enhance node features for training Message Passing Neural Networks (MPNNs) (Gilmer et al., 2017) to predict node categories. Chen et al. (2023) extend the work of  He et al. (2023) by using LLMs both as feature enhancers and as predictors for node classification. They offer several observations such as Chain-of-thoughts is not contributing to performance gains. Guo et al. (2023) perform an empirical study on using LLMs to solve structure and semantic understanding tasks. More recently, Ye et al. (2023) propose InstructGLM for the instruction tuning of LLMs, like LLaMA (Touvron et al., 2023), for node classification tasks.  Zhao et al. (2023) introduces GraphText to encode graph into natural languages using a graph-syntax tree. Our study differs with this line of research in terms of the motivation: while we are using text-rich graph datasets as a testbed, our primary goal is to gain deeper understanding of LLMs’ capability of processing the graph modality, instead of leveraging LLMs to better solve node classification tasks.
@@ 3.1 Research Questions
In this section, we aim to gain a deeper understanding of three central questions. First, does the classification performance of LLMs on common node classification datasets come from data leakage? Second, do LLMs understand the prompts as graph structures? Third, what factors contribute to LLMs benefiting from structural information? To ground our study, we experiment with the ChatGPT API and LLaMA-2-7B model on node classification datasets that have textual node features.
@@ 3.2 Data Leakage as a Potential Contributor of Performance
While LLMs have achieved decent classification performance on common node classification datasets, there is a risk that the performance of LLMs is artificially inflated by data leakage (He et al., 2023; Guo et al., 2023; Hu et al., 2023). Note that most node classification benchmark datasets have a data cut-off at 2019 (see Table 9 in Appendix B.1), but ChatGPT (the model used throughout the paper is gpt-3.5-turbo-0613) was trained on data up to September 2021 (OpenAI, 2023) and LLaMA-2 was trained on data up to September 2022  [^2] https://github.com/facebookresearch/llama/blob/main/MODEL_CARD.md. While the training datasets of ChatGPT and LLaMA-2 are not publicly available, given the widespread of these datasets on the internet and the enormous training corpus of ChatGPT and LLaMA-2, it is reasonable to worry about the data leakage issue on these datasets.
@@
3.3 Are LLMs Treating the Prompts as Graphs or Paragraphs on Text-attributed Graphs
Previous works design templates to encode the ego-graph around the target node into natural languages by listing neighboring nodes’ text information (Chen et al., 2023; Hu et al., 2023) or giving edge indexes (Tang et al., 2023), adjacency lists (Das et al., 2023) or neighborhood summary (Guo et al., 2023). It is implied that some previous works hold that LLMs may benefit from describing the explicit graph structures. However, it is unclear whether LLMs actually understand the prompt as intended. Thus, we ask the question: do LLMs treat the input prompt as a graph, or merely as a paragraph of text with augmented keywords? This question is fundamental in understanding the performance gain after incorporating structural information. If LLMs just process the prompt as a paragraph consisting of neighboring papers’ titles, then the performance gain can only come from the extra phrases included in neighboring nodes, instead of the topological information. We investigate this question by two adversarial experiments:
@@ 3.3 Are LLMs Treating the Prompts as Graphs or Paragraphs on Text-attributed Graphs
To verify whether prompts are processed as linearized paragraphs, we create a linearized version of $k$-hop title, $k$-hop title+label and $k$-hop attention by stripping away descriptive text about the ego-graph structure from the original prompt. For instance, in the original $k$-hop title, a typical introduction like “It has the following neighbor papers at hop $k$:” is removed. In the linearized version, we condense this to a single paragraph comprising all neighbors’ titles, as shown in the last row of Table 1. We then compare the performance of the graph-structured and linearized prompts on ChatGPT. As indicated in Table 4, the results reveal minimal performance disparity across all five prompt styles after linearization. A notable finding is that in the $1$-hop attention category, even the elimination of text specifying neighbors’ importance ranking (“It has the following important neighbors, from most related to least related”) has little impact. These findings imply that LLMs tend to process prompts in a linearized paragraph format.
@@
In summary, our study suggests that LLMs interpret inputs more as contextual paragraphs than as graphs with topological structures. By two adversarial experiments, we show that neither linearizing nor rewiring ego-graph has significant impact on the classification performance of LLMs.
@@ 3.4 Impact of Homophily on LLMs Classification Accuracy
We design a controlled experiment to demonstrate the effect of local homophily ratio on prediction accuracy. We gradually drop neighbors in three different ways: a) drop the neighbors with same label as the target node; b) drop the neighbors with different label as the target node; and c) drop neighbors randomly. We include details about the neighbor dropping strategies in Appendix E. The experimental results are shown in Figure 3, where we observe an evident trend: as we selectively remove neighbors sharing the same labels, there’s a decrease in prediction accuracy. Conversely, discarding neighbors with different labels leads to an increase in accuracy. This selective dropping inherently modifies the local homophily ratio within the prompts. The results show that accuracy of predictions made by LLMs is positively related to local homophily ratio. A similar experiment adding neighbors instead of dropping neighbors in Appendix E also shows a similar trend. Combine this observation with the premise that LLMs actually understand the prompts as a paragraph, we propose that the significance of homophily is intrinsic. This is because LLMs benefit from structural information only when the neighborhood is homophilous, meaning the neighbors contain phrases related to the groundtruth label of the target node. Conversely, diminishing the local homophily ratio by excluding such neighbors can adversely affect prediction accuracy.
@@
For the other datasets, a positive correlation is evident in both the zero-shot and 1-hop title+label settings. In Table 6, zero-shot prediction (the one that doesn’t use structural information at all) also showed high correlation with the homophily ratio of the node. This suggests a complicated mechanism for LLMs to perform better on homophilous nodes: those nodes are easier to be classified in the first place; the added structural information has some additional contributions. Homophilous nodes are easier to classify potentially because nodes that are not homophilous often blend various topics, which makes predicting their category more challenging than homophilous nodes.
@@ 4 Conclusions and Future Work
In conclusion, our study provides key insights into the application of LLMs in processing structured data, particularly in node classification tasks on text-rich graphs. In this study, we have adapted node classification datasets with textual features from graph learning benchmarks to establish a testbed for LLMs augmented with structured data. By curating a new dataset, we show that data leakage is not the major contributor of LLMs on common node classification benchmarks. Our findings also raise critical concerns regarding the actual depth of understanding that LLMs have of structured data. By two adversarial experiments, we show that LLMs actually do not understand the prompts as graphs structures. On the contrary, LLMs understand the prompts more as paragraphs with augmented keywords. As a consequence, LLMs only show improvement from structural information when the neighborhood is homophilous. Moreover, our findings suggest that when the target node itself contains a wealth of relevant phrases, the additional structural information becomes redundant.
@@ Appendix A Details about Prompting Format and Settings
ogbn-arxiv, arxiv-2023 Please predict the most appropriate arXiv Computer Science (CS) sub-category for the paper. The predicted sub-category should be in the format ’cs.XX’.
@@ D.1 Classification accuracy on LLaMA-2-7B-chat
Appendix E Additional Analysis for Dropping experiments
@@
Adding Neighbors instead of Dropping.
Figure 6: Performance comparison of adding neighbors using different strategies across arxiv-2023, cora, and ogbn-product datasets. Three adding strategies are evaluated: “add same” adds neighbors with the same label as the target node; “add different” adds neighbors with different labels as the target node; and “add random” randomly selects neighbors for addition. When percentage is $1$, “add same” strategy adds all same-label neighbors but excludes all different-label neighbors, and “add different” strategy adds all different-label neighbors but excludes all same-label neighbors.
To further investigate whether LLMs are benefiting from structural information, we conducted an additional experiment of adding neighbors instead of dropping neighbors. One could argue that adding neighbors that has different groundtruth labels as the target node can still provide some benefits. The results is shown in 6. For all three datasets, adding neighbors with different labels will decrease the prediction performance.
Appendix F Additional Analysis for Data Leakage
@@
System Prompt Zero-shot 1-hop title+label
@@
Please predict the most appropriate arXiv Computer Science (CS) sub-category for the paper. The predicted sub-category should be in the format ’cs.XX’. 74.0 73.7 74.3 70.4
Please predict the most appropriate arXiv Computer Science (CS) sub-category for the paper. Your answer should be chosen from cs.AI, ..cs.SY. The predicted sub-category should be in the format ’cs.XX’. 66.0 68.1 70.7 67.9
Please predict the most appropriate original arXiv identifier for the paper. The predicted arxiv identifier should be in the format ’arxiv cs.xx’. 71.3 70.8 73.7 67.5
Please predict the most appropriate original arXiv identifier for the paper. Your answer should be chosen from cs.ai,.. cs.sy. The predicted arxiv identifier should be in the format ’arxiv cs.xx’. 58.4 57.2 71.7 64.2
Please predict the most appropriate category for the paper. Your answer should be chosen from "Artificial Intelligence",.. "Systems and Control". 54.6 53.4 74.1 67.8